OpenStudy (anonymous):

a toy box measures 8 inches by 2.6 inches by 5 inches. what is the volume of the toy box?

OpenStudy (diyadiya):

\[Volume = Length \times width \times height \]

OpenStudy (anonymous):

multiply all the numbers

OpenStudy (anonymous):

What exponent?

OpenStudy (diyadiya):

exponent?

OpenStudy (anonymous):

yes, volume is like the answer and then their is a little number at the top, like squared is a 2, volume is a 3 i think

OpenStudy (diyadiya):

\[Volume = (8 \times 2.6 \times 5)inch^3\]